What is the mole fraction of NaCl in a solution of 58.5g NaCl in 180mL H2O?

a. 0.10
b. 0.90
c. 0.091
d. 0.20
1
Expert's answer
2020-04-18T06:59:35-0400

molar mass of NaCl= 23+35.5= 58.5 g/mole

and molar mass of H2O= 18 g/mole

so, here given mass of NaCl= 58.5 and mass H2O = 180 g

Now,

Mole of NaCl="n_1=\frac{mass}{molar mass}=\frac{58.5}{58.5}=1"


mole of H2O="n_2=\frac{mass}{molar mass}=\frac{180}{18}=10"


mole fraction of NaCl="x_1= \frac{n_1}{n_1+n_2}= \frac{1}{1+10}=0.091"